OpenStudy (anonymous):

PLEASE HELP! Use completing the square to change the form of the function. f(x) = 2x^2 + 12x + 17 (If completing the square doesn't work, the quadratic formula may work)

OpenStudy (anonymous):

Use quadratic formula: a=2 b=12 c=17 and get −3±0.5√ 2

OpenStudy (anonymous):

it said this answer was wrong

OpenStudy (anonymous):

I got ± √2/2 - 3 and it said that was wrong too

OpenStudy (anonymous):

to change the form of the function. f(x) = 2x^2 + 12x + 17 you might arrived at the form f(x)=(x +____ ?)(x +______? )

OpenStudy (anonymous):

I'm not sure...

OpenStudy (anonymous):

I only have 1 h0ur left! please help me!!

OpenStudy (anonymous):

any suggestions? the 2 solutions posted are somehow wrong.

OpenStudy (anonymous):

-6 plus or minus square root of 2 all divided by 4

OpenStudy (anonymous):

Is this assuming 2x^2 + 12x + 17 equals zero? @sm101

OpenStudy (anonymous):

Thats something I wasn't sure of @nicetomeetyou , thats exactly how the problem is written, but I assume it may be = 0 otherwise how could you solve it?

OpenStudy (anonymous):

I mean the porblem is exactly written as ... Use completing the square to change the form of the function. f(x) = 2x^2 + 12x + 17

OpenStudy (anonymous):



OpenStudy (anonymous):

Ok! just 1 sec...

OpenStudy (anonymous):

f(x) = 2x^2 + 12x + 17 dive this by 2 we get x^2+6x+ 17/2 ..now on that 6x, 6/2=3 therefore use (x+3)^2 +17/2 -9 =(x+3)^2 -1/2 ...yes or no?